# Speech to Speech Translation with Translatotron: A State of the Art Review

### v1

@@ 6 Comparative study
The first line of Table 3, lists all the models and the features used for the comparison. On line 2 the feature considered is the architecture with Basic Seq to Seq for Translatotron 1, improved Seq to Seq for Translatotron 2, and Advanced Transformer for Translatotron 3. On the last line, the feature considered is the sampling rate: with 16 kHz for translatotron 1, and 2 and a range of values from 16 kHz to 48 kHz for translatotron 3. This table clearly shows that translatotron 3 is the best option when designing a sequence-to-sequence S2ST model. When designing an English to Yoruba translation model with a health environment as a context, we believe translatotron 3 to be the best option.

### v2

@@ 6 Comparative study
The first line of Table 3, lists all the models and the features used for the comparison. On line 2 the feature considered is the architecture with Basic Seq to Seq for Translatotron 1, improved Seq to Seq for Translatotron 2, and Advanced Transformer for Translatotron 3. On the last line, the feature considered is the sampling rate: with 16 kHz for translatotron 1, and 2 and a range of values from 16 kHz to 48 kHz for translatotron 3. This table clearly shows that translatotron 3 is the best option when designing a sequence-to-sequence S2ST model. When designing an English-to-Yoruba translation model with a health environment as a context, we believe translatotron 3 to be the best option.
Figure 4 shows the evolution of the speech-to-speech translation models. The tree shows that everything started in 1990 with the cascade models. The first sequence-to-sequence model appears between 2012 and 2014. The latest version of the speech-to-speech translation the translatotron 3 model was released in 2024.
Figure 4: Speech-to-Speech translation Evolution tree
@@ Conclusion
Acknowledgement
The authors acknowledge Google for providing funding to the second author through the 2024 Google Academic Research Award for this study. The authors also acknowledge the Covenant Applied Informatics and Communication Africa Centre of Excellence (CApICACE) for providing funding towards the publication of this work.